OpenStudy (anonymous):

Help please!! It's an infinite serie \[\displaystyle\sum_{n=1}^\infty \dfrac{n}{3^n}\]

OpenStudy (anonymous):

\[\displaystyle\sum_{n=1}^\infty \dfrac{n}{3^n} \]

OpenStudy (anonymous):

please i need the solution

OpenStudy (anonymous):

i think it is the derivative of a geometric series

OpenStudy (anonymous):

without the \(n\) up top it would be \(\frac{1}{1-\frac{1}{3}}\) but with the \(n\) it is \[\frac{1}{(1-\frac{1}{3})^2}\]

OpenStudy (anonymous):

oh and if you are starting at \(n=1\) instead of \(n=0\) you have to subract \(1\)

OpenStudy (anonymous):

that is not right what i just said since you are starting at \(n=1\) it is \[\frac{\frac{1}{3}}{(1-\frac{1}{3})^2}\]

OpenStudy (anonymous):